How do i do this question. A weightlifter lifts a 32 kg set of weights from ground level to position over the head. The vertical lift is 2.2 m. How much work does the weightlifter do
1
Expert's answer
2013-03-04T08:50:01-0500
Let's write the law conservation of energy(the work which the weightlifter do will equal to difference between initial and final potential energies):
A = U2 - U1
A - work which the weightlifter do,
U1, U2 - initial and final potential energies
U = mgh - the definition of the potential energy
m - mass
g - gravitational acceleration
h - high
At the beginning, the weightlifter was staying on the ground, so h=0.
A = mgh - mg*0 = mgh
A = 32*2.2*9.8 = 689.92 J
Answer: 689.92 J